# Irregular Tensor Low-Rank Representation for Hyperspectral Image Representation

### v1

@@ Abstract
Spectral variation is a common problem for hyperspectral image (HSI) representation. Low-rank tensor representation is an important approach to alleviate spectral variations. However, the spatial distribution of the HSI is always irregular, while the previous tensor low-rank representation methods can only be applied to the regular data cubes, which limits the performance. To remedy this issue, in this paper we propose a novel irregular tensor low-rank representation model. We first segment the HSI data into several irregular homogeneous regions. Then, we propose a novel irregular tensor low-rank representation method that can efficiently model the irregular 3D cubes. We further use a non-convex nuclear norm to pursue the low-rankness and introduce a negative global low-rank term that improves global consistency. This proposed model is finally formulated as a convex-concave optimization problem and solved by alternative augmented Lagrangian method. Through experiments on four public datasets, the proposed method outperforms the existing low-rank based HSI methods significantly. Code is available at: https://github.com/hb-studying/ITLRR.
@@
I INTRODUCION
With the development of hyperspectral remote sensing technology, hyperspectral image (HSI) can capture spectral information across hundreds of contiguous spectral bands, as illustrated in Fig. 1a. This rich band information enables discrimination among various materials and facilitates applications in agriculture [agricultural], environmental monitoring [environment_monitor2] and urban planning [urban_planning]. As HSIs often encompass a limited number of materials and there are high correlations among the spectral signatures of these materials, HSIs inherently exhibit a low-rank structure [structure_of_HSI]. However, owing to sensor interference or variations in imaging conditions, the spectral signatures of pixels within homogeneous regions may vary dramatically [2014HRSI]. Such spectral variations will disrupt the low-rank structure significantly.
Therefore, many methods based on the low-rank property of HSI have been proposed to alleviate the spectral variation and enhance representation ability. Several methods [2016HRSI-RPCA], [2018FRPCALG] are based on Robust Principal Component Analysis (RPCA [2011rpca]) that unfold three-dimensional hyperspectral images into matrices, assuming that the matrix can be decomposed into a low-rank matrix and a sparse noise matrix. Later, considering that similar materials usually appear in local regions, as illustrated in Fig. 1a, many methods [2014ILRA], [2015-adjust-LRMR], [2017SSLR], [2018S3LRR] partition the HSI into rectangular patches as illustrated in Fig. 1b and restore each patch individually using RPCA-based techniques. To capture local information more effectively, superpixel segmentation algorithms have been employed to obtain irregular homogeneous regions in [2017SSLRR], [2018SLRA_GBM], [2017Multiscale_SLRA] as illustrated in Fig. 1c.
Nonetheless, it is essential to recognize that the data structure of HSI is intrinsically tensor rather than matrix [2014HRSI]. The above methods transform HSI data into the matrix form, potentially corrupting the spatial features [2020TRPCA]. In contrast, tensor representations can retain spatial information effectively and yield superior restoration results. Tensor-based robust principal component analysis (TRPCA) models decompose the input HSI tensor into low-rank and sparse tensors, which are widely applied in HSI [trpca_13], [trpca_14], [trpca_15], [trpca_11], [2020LSSTRPCA], [2022LPGTRPCA]. However, it is important to note that the above tensor-based methods can only handle the regular data cubes. While as shown in Fig. 1, the same material in HSI is usually distributed locally in irregular cubes, the global tensor low-rank representation on the whole HSI cube cannot effectively exploit the irregular data distribution.
@@ I INTRODUCION
To this end, we propose a novel Irregular Tensor Low-Rank Representation (ITLRR), which can process the irregular three-dimensional data cubs. Fig. 2 visually illustrates the proposed method. First, we employ an off-the-shelf superpixel segmentation technique [ERS] to split the hyperspectral data into irregular 3D blocks. Second, we complete each irregular 3D data cube by using blank tensors to make up a regular data cube, which may reflect the characteristic of the irregular 3D cube. Then we apply tensor nuclear norm to the completed 3D cubes, and we design a novel operation where only the irregular cubes make contribution to the nuclear norm, while the black cubes are hided in LRR. This operation will empower the TLRR to model the irregular 3D cubes. We further design a non-convex tensor nuclear norm that closely approximates the true rank. However, over-emphasizing the low-rankness in the local patch may lead to the over-smoothing representation, i.e., the representation of the non-dominated class will be close to that of the dominated class in a local patch. Therefore, we propose a negative low-rank term to enhance discriminative ability of the low-rank representation. The proposed model is formulated as a convex-concave problem and solved by alternative augmented Lagrangian method. Extensive experimental results demonstrate that our method surpasses state-of-the-art low-rank based methods significantly. The contributions of this study can be summarized as follows:
@@
    Our proposed method is the first to explore the low-rank properties of the irregular tensors.
@@
The rest of this paper is organized as follows. Section II briefly introduces the notations and preliminaries used in the paper and reviews existing methods based low-rank representation for HSI. After that, the proposed ITLLR method is presented in Section III. In Section IV, we demonstrate experimental evaluations and comparisons, followed by the conclusions in Section V.
Figure 2: Framework of the proposed method. (I) stands for the origin HSI with spectral variation. The same color indicates that those pixels belong to the same class, and the black dot represents the data noise. We first divide the input HSI into several irregular 3D cubes by a typical superpixel segmentation method ERS; Then we complete the irregular 3D data cubes with black shape-complementary data cubes to constitute regular tensor patches (III), and design a novel ITLRR model (IV) that can pursue the tensor low-rank representation, which only relies on the original irregular data cubes and ignores the black complementary 3D patches. Besides, the ITLRR is based on a non-convex tensor nuclear norm that can better approximate the tensor low-rankness (IV). The above irregular tensor low-rank representation process the 3D patches individually, which may lead to the over-smoothing problem, i.e., two disconnected areas with the same material do not share the similar representation (e.g., E1 and E2 in (IV)). To address the problem, we propose to add a negative low-rank term on the whole HSI to enhance the overall discriminative ability (V). Finally, we pack the low-ranked irregular 3D cubes to get the final low-rank HSI representation (VI).
II PRELIMINARIES AND RELATED WORK
HSIs are used for various applications because of the rich band information. However, due to defective imaging conditions, HSI usually suffers from the spectral variation problem, i.e., the same material does not own the same spectral signature. Low-rank representation is an important approach to solve the spectral variation problem. In this section, we first introduce some notations and the preliminaries on tensor low-rank representation, and then review the related low-rank representation methods for HSI.
II-A Notations and Preliminaries
In this paper, we denote scalars by lowercase letters, e.g., $a$, vectors by bold lowercase letters, e.g., $\mathbf{a}$, 2-D matrices by bold uppercase letters, e.g., $\mathbf{A}$, and 3D tensors by calligraphic uppercase letters, e.g., $\mathcal{A}$. And $a_{ijk}$ is the $(i,j,k)_{th}$ entry of $\mathcal{A}$ and $\mathbf{A}^{(k)}$ is the $k$-th frontal slice of $\mathcal{A}$. More notations are summarized in Table I. We will introduce some preliminaries for tensor nuclear norm.
@@ II-A Notations and Preliminaries
1. $\overline{\mathcal{A}}$ is the dicrete Fast Fourier Transform (FFT) of $\mathcal{A}$, i.e.,
@@
3. $\mathscr{F}_{3}\left(\mathbf{A}_{(3)}\right)=\mathcal{A}$, i.e.,
   $\mathbf{A}_{(3)}\in\mathbb{R}^{n_{1}n_{2}\times n_{3}}$ is the third order unfolding of $\mathcal{A}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$.
@@
Let $\mathbf{A}\in\mathbb{R}^{n_{1}\times n_{2}}$ be an matrix, the Singular Value Decomposition (SVD) of $\mathbf{A}$ is a factorization of the form:
@@
[theorem3] Let $\mathbf{A}\in\mathbb{R}^{n_{1}\times n_{2}}$ be an arbitrary matrix. The partial of $\|\mathbf{A}\|_{*}$ is $\partial\|\mathbf{A}\|_{*}=\{\mathbf{UV^{T}}+\mathbf{Z}:\mathbf{Z}\in\mathbb{R}^{n_{1}\times n_{2}},\mathbf{U^{T}Z}=0,\mathbf{ZV}=0,\|\mathbf{Z}\|_{2}\leq 1\}$, where $\mathbf{A}=\mathbf{USV^{T}}$ and $\|\mathbf{Z}\|_{2}=\sqrt{\sum^{n_{1}}_{i}\sum^{n_{2}}_{j}a_{ij}^{2}}$.
@@
[theorem4] For any $\lambda>0$ and $\mathbf{Y}\in\mathbb{R}^{n_{1}\times n_{2}}$, a globally optimal solution to the following optimization problem
@@
For a tensor $\mathcal{A}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$, its nuclear norm defined in [2020TRPCA] is
@@ II-B Related Work
Due to the inherent low-rank property of HSIs, many low-rank-based methods have been proposed to remedy the spectral variation problem. For example, 2016HRSI-RPCA reformatted HSI data into a two-dimensional matrix and applied RPCA to capture the low-rank representation that effectively mitigates spectral variations. Further, 2018FRPCALG assumed that HSIs exhibit a clustering property in the unified subspace and incorporated a Laplacian graph term into the RPCA model to generate the low-rank matrix.
@@
The above methods do not reveal the local spatial correlation of pixel points in the same region. Therefore, many methods were proposed to exploit this prior. For example, 2014ILRA divided the HSIs into patches and restored each patch separately by using Go Decomposition algorithm [2011GoDec]. Considering the variability of noise intensity across different bands, 2015-adjust-LRMR introduced an iterative framework with noise adjustments employing patch-wise low-rank approximation method for HSI. 2017SSLR employed RPCA on patches to explore local spatial information and applied RPCA individually on each spectral band to investigate spatial information. Moreover, 2018S3LRR proposed a model exploring the low-rank property spectral and spatial domain simultaneously. Because homogeneous regions are always irregular, 2017SSLRR, 2018SLRA_GBM introduced superpixel segmentation to preserve spatial correlation of pixels in homogeneous areas. 2017Multiscale_SLRA proposed a strategy of multiscale superpixels to solve the problem of determining the optimal superpixel size.
@@
The above matrix-based methods may corrupt spatial features due to the intrinsic structure of HSI is a 3D tensor. To address this issue, many tensor-based methods have been proposed. tubal_rank, and 2014TRPCA designed the tensor multi-rank and the tensor tubal rank respectively. 2020TRPCA defined a new tensor nuclear norm which is consistent with the matrix nuclear norm and proposed a TRPCA model which decomposes the 3D HSI tensor into a low-rank tensor and a error tensor. trpca_13 introduced the TRPCA into anomaly detection and proposed a low-rank tensor decomposition based anomaly detection algorithm. trpca_14 proposed a simultaneous reconstruction and anomaly detection framework for compressive hyperspectral imaging based on TRPCA model. trpca_15 considered both the outliers and different types of noise and proposed an HSI denoising model based on the robust low-rank tensor recovery. trpca_11 incorporated spectral graph regularization into TRPCA and proposed the graph-regularized TRPCA algorithm for HSI denoising. 2020LSSTRPCA proposed a lateral-slice sparse tensor RPCA with a tensor $l_{2,1}$ norm for sparse component to gross errors or outliers. 2022LPGTRPCA incorporated a tensor $l_{2,2,1}$ norm for frontal slice sparsity and a position-based Laplacian graph for local structure preservation to improve hyperspectral image classification performance.
II-C Motivation
While tensor-based methods can effectively preserve the spatial structure of the HSIs, they can only be applied to the regular 3D data cubes, overlooking the fact that the same material in an HSI is usually distributed within irregular homogeneous regions. To solve this problem, we propose a novel tensor low-rank representation model that can process irregular data cubes and accordingly promote the representation ability.

### v3

@@ Abstract
Spectral variations pose a common challenge in analyzing hyperspectral images (HSI). To address this, low-rank tensor representation has emerged as a robust strategy, leveraging inherent correlations within HSI data. However, the spatial distribution of ground objects in HSIs is inherently irregular, existing naturally in tensor format, with numerous class-specific regions manifesting as irregular tensors. Current low-rank representation techniques are designed for regular tensor structures and overlook this fundamental irregularity in real-world HSIs, leading to performance limitations. To tackle this issue, we propose a novel model for irregular tensor low-rank representation tailored to efficiently model irregular 3D cubes. By incorporating a non-convex nuclear norm to promote low-rankness and integrating a global negative low-rank term to enhance the discriminative ability, our proposed model is formulated as a constrained optimization problem and solved using an alternating augmented Lagrangian method. Experimental validation conducted on four public datasets demonstrates the superior performance of our method compared to existing state-of-the-art approaches. The code is publicly available at https://github.com/hb-studying/ITLRR.
@@
I Introduction
With the advancement of hyperspectral remote sensing technology, hyperspectral images (HSIs) can capture spectral data across hundreds of contiguous bands, as shown in Fig. 1. This rich band information enables discrimination among various materials and facilitates applications in agriculture [1], environmental monitoring [2] and urban planning [3]. Since HSIs typically contain a limited set of materials and exhibit high correlation among their spectral signatures, they inherently possess a low-rank structure [4]. However, due to sensor interference or variations in imaging conditions, the spectral signatures of pixels in homogeneous regions may vary considerably [5], disrupting this low-rank structure.
To address spectral variation and enhance representation ability, numerous methods leveraging the low-rank property of HSIs have been proposed. Several methods [6, 7] are based on Robust Principal Component Analysis (RPCA [8]) that unfold three-dimensional hyperspectral images into matrices, assuming that the matrix can be decomposed into a low-rank matrix and a sparse noise matrix. Later, considering that similar materials usually appear in local regions, as illustrated in Fig. 1, many methods [9, 10, 11, 12, 13] partition the HSI into rectangular patches as illustrated in Fig. 1 and restore each patch individually using RPCA-based techniques. To capture local information more effectively, superpixel segmentation algorithms have been employed to obtain irregular homogeneous regions in [14, 15, 16, 17] as illustrated in Fig. 1.
Nonetheless, it is essential to recognize that the data structure of HSI is intrinsically tensor-based, not matrix-based [5]. The above methods transform HSI data into matrix form, potentially corrupting the spatial features [18]. In contrast, tensor representations can preserve spatial information effectively and yield superior restoration results. Tensor-based robust principal component analysis (TRPCA) models decompose the HSI tensor into low-rank and sparse components, which are widely applied to HSI [19, 20, 21, 22, 23, 24]. However, it is important to note that the above tensor-based methods can only handle regular data cubes. While as shown in Fig. 1, the same material in HSI is typically distributed in irregular, localized regions, the global tensor low-rank representation applied to the entire HSI cube is not well-suited for capturing the irregular data distribution.
@@ I Introduction
To this end, as shown in Fig. 2, we propose a novel Irregular Tensor Low-Rank Representation (ITLRR), which can effectively process the irregular three-dimensional data cubs. For the irregular 3D data, we first complete it by using blank tensors to make up a regular data cube, and then we apply distinct low-rank and sparse constraints to the complementary data cubes and the original irregular data cube, which enables the effective exploration of the low-rank representation of irregular 3D data cube. We further design a non-convex tensor nuclear norm that closely approximates the true rank, and propose a global negative low-rank term to enhance discriminative ability of the overall low-rank representation. The proposed model is formulated as a optimization problem and solved by alternative augmented Lagrangian method. Extensive experimental results demonstrate that our method surpasses state-of-the-art low-rank based and deep-learning based methods significantly. The main contributions of this study are summarized as follows:
@@
    Our proposed method is the first to explore the discriminative low-rank properties of the irregular tensors.
@@
The rest of this paper is organized as follows. Section II introduces the notations and preliminaries used in this paper, along with a review of existing methods for HSI. The proposed ITLRR method is presented in Section III. Section IV provides experimental evaluations and comparisons, followed by the conclusions in Section V.
Figure 2: Framework of the proposed method. (I) stands for the origin HSI with spectral variation. The same color indicates that those pixels belong to the same class, and the black dot represents the data noise. We first divide the input HSI into several irregular 3D cubes by a typical superpixel segmentation method ERS; Then we complete the irregular 3D data cubes with black shape-complementary data cubes to constitute regular tensor patches (III), and design a novel ITLRR model (IV) that can pursue the tensor low-rank representation, which only relies on the original irregular data cubes and ignores the black complementary 3D patches. Besides, the ITLRR is based on a non-convex tensor nuclear norm that can better approximate the tensor low-rankness (IV). The above irregular tensor low-rank representation processes the 3D patches individually, which may lead to the over-smoothing problem, i.e., two disconnected areas with the same material do not share a similar representation (e.g., E1 and E2 in (IV)). To address the problem, we propose to add a negative low-rank term on the whole HSI to enhance the overall discriminative ability (V). Finally, we pack the low-ranked irregular 3D cubes to get the final low-rank HSI representation (VI).
@@
1. $\overline{\mathcal{A}}$ is the discrete Fast Fourier Transform (FFT) of $\mathcal{A}$, i.e.,
@@
3. $\mathbf{A}_{(3)}\in\mathbb{R}^{n_{1}n_{2}\times n_{3}}$ is the third order unfolding of $\mathcal{A}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$,
   which can be expressed as $\mathscr{F}_{3}\left(\mathbf{A}_{(3)}\right)=\mathcal{A}$.
@@
II Preliminary and Related Work
HSIs are widely used in various applications due to their rich spectral information. However, due to imperfect imaging conditions, HSIs often suffer from the spectral variation problem, where the same material may exhibit different spectral signatures. Low-rank representation is an effective approach to address this issue. In this section, we first introduce the notations and preliminaries related to tensor low-rank representation, followed by a review of existing methods for HSIs.
II-A Notations and Preliminaries
In this paper, we denote scalars by lowercase letters, e.g., $a$, vectors by bold lowercase letters, e.g., $\mathbf{a}$, 2-D matrices by bold uppercase letters, e.g., $\mathbf{A}$, and 3D tensors by calligraphic uppercase letters, e.g., $\mathcal{A}$. And $a_{ijk}$ is the $(i,j,k)_{th}$ entry of $\mathcal{A}$ and $\mathbf{A}^{(k)}$ is the $k$-th frontal slice of $\mathcal{A}$. More notations are summarized in Table I. We will introduce some preliminaries for the tensor nuclear norm.
@@ II-A Notations and Preliminaries
Let $\mathbf{A}\in\mathbb{R}^{n_{1}\times n_{2}}$ be a matrix, the Singular Value Decomposition (SVD) of $\mathbf{A}$ is a factorization of the form:
@@
[25] Let $\mathbf{A}\in\mathbb{R}^{n_{1}\times n_{2}}$ be an arbitrary matrix. The partial derivative of $\|\mathbf{A}\|_{*}$ is $\partial\|\mathbf{A}\|_{*}=\{\mathbf{UV^{T}}+\mathbf{Z}:\mathbf{Z}\in\mathbb{R}^{n_{1}\times n_{2}},\mathbf{U^{T}Z}=0,\mathbf{ZV}=0,\|\mathbf{Z}\|_{2}\leq 1\}$, where $\mathbf{A}=\mathbf{USV^{T}}$ and $\|\mathbf{Z}\|_{2}=\sqrt{\sum^{n_{1}}_{i}\sum^{n_{2}}_{j}a_{ij}^{2}}$.
@@
[26] For any $\lambda>0$ and $\mathbf{Y}\in\mathbb{R}^{n_{1}\times n_{2}}$, a globally optimal solution to the following optimization problem
@@
For a tensor $\mathcal{A}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$, its nuclear norm defined in [18] is
@@
Definition 2 (The $\oplus$ operator).
For a regular tensor $\mathcal{X}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$, composed of several non-overlapping irregular tensors $\mathcal{X}_{1},\mathcal{X}_{2},\dots,\mathcal{X}_{k}$, it can be expressed as $\mathcal{X}=\mathcal{X}_{1}\oplus\mathcal{X}_{2}\oplus\cdots\oplus\mathcal{X}_{k}$. Specially, the regular tensor is indexed by a matrix $\mathcal{I}\in\mathbb{R}^{n_{1}\times n_{2}}$. Each irregular tensor $\mathcal{X}_{i}$ is associated with a unique index set $\mathcal{I}_{i}$, i.e, $\mathcal{I}_{i}\cap\mathcal{I}_{j}=\emptyset,\forall i\neq j$ and the union of all index sets satisfies $\bigcup_{i=1}^{k}\mathcal{I}_{i}=\mathcal{I}$. Each index set $\mathcal{I}_{i}$ determines the positions of the $i$-th irregular tensor $\mathcal{X}_{i}$ in the regular tensor $\mathcal{X}$, i.e, $\mathcal{X}(a,b,:)=\mathcal{X}_{i}(a,b,:),\forall(a,b)\in\mathcal{I}_{i}$. The $\oplus$ operator is then defined to merge these irregular tensors based on their index mappings:
$\mathcal{X}_{i}\oplus\mathcal{X}_{j}=\mathcal{X}(a,b,:),\forall(a,b)\in\mathcal{I}_{i}\cup\mathcal{I}_{j}.$ (5)
For example, for a regular tensor $\mathcal{X}\in\mathbb{R}^{8\times 4\times h}$ composed of three irregular tensors, the process of merging based on the $\oplus$ operation is illustrated in Fig. 3.
Figure 3: The illustration of the operation $\oplus$ to combine three irregular 3D patches.
@@ II-B Related Work
Due to the inherent low-rank property of HSIs, many low-rank-based methods have been proposed to remedy the spectral variation problem. For instance, Mei et al. [6] restructured HSI data into a two-dimensional matrix and applied RPCA to capture the low-rank representation, effectively reducing spectral variations. Additionally, Sun and Du [7] assumed that HSIs exhibit a clustering property within a unified subspace and incorporated a Laplacian graph term into the RPCA model to enhance the low-rank matrix generation.
@@
The methods mentioned above do not reveal the local spatial correlation of pixel points within the same region. To address this, several approaches have been proposed to leverage this spatial prior. For example, Zhang et al. [9] divided the HSIs into patches and restored each patch separately using the Go Decomposition algorithm [27]. Considering the variability of noise intensity across different bands, He et al. [10] introduced an iterative framework with noise adjustments employing a patch-wise low-rank approximation method for HSI. Mei et al. [11] employed RPCA on patches to explore local spatial information and applied RPCA individually on each spectral band to investigate spatial information. Moreover, Mei et al. [12] proposed a model exploring the low-rank property spectral and spatial domain simultaneously. Since homogeneous regions are often irregular, Fan et al. [14] and Mei et al. [15] incorporated superpixel segmentation to preserve the spatial correlation of pixels in homogeneous regions. Zhang et al. [16] proposed a strategy of multiscale superpixels to solve the problem of determining the optimal superpixel size.
@@
The above matrix-based methods may corrupt spatial features due to the inherent 3D structure of HSIs. To address this, many tensor-based methods have been proposed. Kilmer and Martin [28], and Zhang et al. [29] designed the tensor multi-rank and the tensor tubal rank respectively. Lu et al. [18] defined a new tensor nuclear norm consistent with the matrix nuclear norm and proposed a TRPCA model that decomposes the 3D HSI tensor into a low-rank tensor and an error tensor. Li et al. [21] considered both the outliers and different types of noise and proposed an HSI denoising model based on the robust low-rank tensor recovery. Nie et al. [22] incorporated spectral graph regularization into TRPCA and proposed the graph-regularized TRPCA algorithm for HSI denoising. Sun et al. [23] proposed a lateral-slice sparse tensor RPCA with a tensor $l_{2,1}$ norm to model sparse components and handle gross errors or outliers. Wang et al. [24] incorporated a tensor $l_{2,2,1}$ norm for frontal slice sparsity and a position-based Laplacian graph to preserve the local structure, improving hyperspectral image classification performance. Xue et al. [30] proposed a tensor convolution-like low-rank dictionary framework that accounts for the shift-invariant low-rankness of tensor data and integrated it into a TRPCA model. Jiang et al. [31] and Zhang et al. [32] considered the irregularity of the real data distribution; however, they overlooked the discriminability of the representation.
While tensor SVD-based methods have shown promising results in capturing the low-rank structure of HSIs, other tensor decomposition techniques, such as Tucker decomposition and Tensor Ring (TR) decomposition, have also gained significant attention in recent years [33]. For instance, Xu et al. [34] introduced a hyperspectral image super-resolution framework based on Tucker decomposition, incorporating $l_{1}$-norm regularization on the core tensor to model sparsity and unidirectional total variation on dictionary matrices to ensure piecewise smoothness. Similarly, Tian et al. [35] developed a mixed noise removal model for HSIs by combining low-rank Tucker decomposition with $l_{0}$-norm-based regularizers, effectively capturing global correlations and improving sparse noise removal performance. In the context of TR decomposition, Xu et al. [36] proposed a high-order coupled TR representation model for hyperspectral image super-resolution, which leverages shared latent core tensors and graph-Laplacian regularization to preserve spectral information. Zhang et al. [37] designed a hyperspectral image fusion model based on TR decomposition, integrating logarithmic tensor nuclear norm regularization and weighted total variation to enhance spatial-spectral continuity. Wan et al. [38] proposed a subspace-based hyperspectral image reconstruction model that utilizes spectral quadratic variation regularized autoweighted TR decomposition.
With the rapid advancement of deep learning, it has been widely applied in HSI processing. Convolutional neural networks (CNNs) have been utilized to extract spatial contextual information from HSIs using 2D or 3D convolutional kernels [39, 40, 41, 42, 43, 44]. For example, Lu et al. [42] introduced a separable deep graph convolutional network that combines CNNs with prototype learning to enhance discriminative feature extraction. More recently, transformers have emerged as a powerful tool for hyperspectral image classification [45, 46, 47, 48, 49]. For instance, Hong et al. [46] proposed a transformer-based method that integrates a group-wise spectral embedding module and a cross-layer adaptive fusion module. Beyond conventional architectures, Li et al. [50] combined low-rank representation with deep learning techniques, enabling automatic parameter learning for hyperspectral anomaly detection. Hong et al. [51] proposed the first spectral remote sensing foundation model using a 3D generative pretrained transformer, incorporating million-scale progressive training and spectral-spatial coupling via tensor masking and multi-target reconstruction.